OpenStudy (anonymous):

You have a large flashlight that takes 4 D-cell batteries, each with a voltage of 1.5 volts. If the current in the flashlight is 2.0 amps, what is the resistance? Select one of the options below as your answer: A. 12 ohms B. 0.75 ohms C. 3.0 ohms D. 0.5 ohms E. 6.0 ohms

OpenStudy (fifciol):

using ohm's law R=V/I = 4*1,5/2= 3 ohms Answer C.

OpenStudy (anonymous):

TY^

OpenStudy (anonymous):

can i draw pic and u answer it its which of the symbols on a circuit does this represent

OpenStudy (anonymous):

A. ammeter B. voltmeter C. switch D. AC power source E. DC power sourc

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry bad drawing it supposed to nbe closer

OpenStudy (fifciol):



OpenStudy (anonymous):

ok cool lol :)